OpenStudy (anonymous):

2x - 55 = 3y 3y = 99 - 9x Using the two equations above, solve for y. A. y= -165 B. y= 14 C. y= 9 D. y= -9

OpenStudy (anonymous):

3y=2x-55 and 3x=99-9x, so 2x-55=99-9x 2x-55=99-9x 2x+9x=99+55 11x=154 x=154/11 x=14 3y=2x-55 3y=2*14-55 3y=28-55 3y=-27 y=-9 The right answer choice is D.

OpenStudy (anonymous):

thanks